OpenStudy (anonymous):

Rationalize the denominator and simplify the quad?root of 2x^5/3w

OpenStudy (anonymous):

\[\sqrt[4]{2x ^{5}/3w}\]

OpenStudy (anonymous):

I came up with \[x ^{4}\sqrt{(54w ^{3}x)/3w}\] but it looks wrong to me

OpenStudy (anonymous):

I'm confused >.<

OpenStudy (anonymous):

ok multiply and divide by 81w^4

OpenStudy (anonymous):

oh sorry...don't do that

OpenStudy (anonymous):

what you need is 81 w^4 inside the root

OpenStudy (anonymous):

so multiply divide 27w^3

OpenStudy (anonymous):

which you did right ?

OpenStudy (anonymous):

if you do that you will get 81w^4 in denominator which when comes out of the root becomes 3w

OpenStudy (anonymous):

Oh ok.. gotchya so far..

OpenStudy (anonymous):

in numerator you will get 54x^5 *w^3

OpenStudy (anonymous):

So multiply divide.. 27w^3 which yields to 81w^4 that becomes 3w out of the root to recap so far..

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So I got the 54, where I messed up is the w and x

OpenStudy (anonymous):

i think it should become x (54 * w^3* x)^1/4 --- 3w

OpenStudy (anonymous):

54 = 2*3*3*3 now you can't have anything else out of the root so the final should be x/3w * ( 54*w^3*x)^1/4

OpenStudy (anonymous):

Ok.. thank you very much for you help with this.. these things are a pain in the butt to me >.<

OpenStudy (anonymous):

lolx.